# Hybrid User Interfaces: Past, Present, and Future of Complementary Cross-Device Interaction in Mixed Reality

### v1

@@ Abstract
We investigate hybrid user interfaces (HUIs), aiming to establish a cohesive understanding and adopt consistent terminology for this nascent research area. HUIs combine heterogeneous devices in complementary roles, leveraging the distinct benefits of each. Our work focuses on cross-device interaction between 2D devices and mixed reality environments, which are particularly compelling, leveraging the familiarity of traditional 2D platforms while providing spatial awareness and immersion. Although such HUIs have been prominently explored in the context of mixed reality by prior work, we still lack a cohesive understanding of the unique design possibilities and challenges of such combinations, resulting in a fragmented research landscape. We conducted a systematic survey and present a taxonomy of HUIs that combine conventional display technology and mixed reality environments. Based on this, we discuss past and current challenges, the evolution of definitions, and prospective opportunities to tie together the past 30 years of research with our vision of future HUIs.
@@
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Immersive augmented reality (AR) and virtual reality (VR) is gradually gaining relevance in everyday life, with affordable off-the-shelf hardware becoming increasingly available to consumers. These mixed reality (MR) platforms [^1] In this work, mixed reality refers to both AR and VR (cf. [136]), as discussed in Sec. 4.3. have now evolved to the point where researchers can explore the nuances of interaction design without being constrained by major technological limitations. In this context, Feiner and Shamash [47] proposed in 1991 the concept of hybrid user interfaces (HUIs), which combine “heterogeneous display and interaction device technologies.” This concept theoretically allows for an infinite integration of technologies, yet, such heterogeneous combinations are especially compelling for MR environments: The ubiquity, convenience, and familiarity of conventional 2D platforms (e.g., smartphones, desktops) provide a perfect complement to the immersion and complexity of optical see-through (OST) and video see-through (VST) head-worn devices (HWDs) [^2] We refer to “head-worn devices” (HWDs) instead of “head-mounted displays” (HMDs) to emphasize the increase in wearability and capabilities of current hardware, but intentionally kept the previous term HMD for describing older hardware.. Yet, although the concept of HUIs has persisted for decades, no coherent delineation has emerged. As a result, there is a distinct lack of consistent design models and terminologies—fragmenting the research community across overlapping research areas such as cross-reality systems [4], transitional interfaces [53], or cross-device interaction [16].
For example, the seminal cross-device taxonomy by Brudy et al. [16] provides an overarching model for the research area of HUIs. However, their taxonomy only considers established devices (e.g., mobile devices) and homogeneous combinations, leaving topics such as MR environments largely for future work. Yet, without firm design principles and in the face of countless possible device combinations, the design space of HUIs can appear bewilderingly large. Therefore, we extend the existing taxonomy by examining complementary cross-device interaction between traditional 2D device technologies (e.g., smartphones, desktops) and novel MR platforms as a prevalent subset of HUIs. By first examining this subset, we take a first step towards a better understanding of the broader research area of HUIs. Thus, we aim to establish a common terminology, allowing researchers and practitioners to better benefit from shared insights, establish consistent frameworks, and inspire future systems. This would, in turn, enable the creation of a cohesive understanding of the unique design possibilities and challenges of HUIs.
@@
  • A positioning of HUIs, creating a better understanding of the term within the current fragmented research landscape (Sec. 2 and Sec. 7).
@@ 2 What is a Hybrid User Interface?
Since 1991, the term “hybrid user interface” has continued to evolve. Despite significant advances in device technologies, no coherent delineation has emerged that delineates HUIs from related research areas such as cross-device interaction. This lack of common understanding makes it difficult to compare results and share insights. In the following, we first review prior definitions and the background (Sec. 2.1) of HUIs, building on previous work by Satkowski and Méndez [125], and position HUIs in relation to current adjacent research areas (Sec. 2.2). Based on this, we discuss how we can identify HUIs (Sec. 2.3) and derive three attributes of HUIs that specifically combine MR environments with conventional display technologies (Sec. 2.4) to guide our literature survey.
Figure 1: We take a look at 30 years of hybrid user interface research and its evolution, focusing on combinations of conventional 2D devices with mixed reality environments. From left to right, (A) starting with its definition [47], (B) early extensions [13] (Image © IEEE 2006), and (C) current usage for augmenting displays [119] (Image © Reipschläger et al., $C$ BY 4.0), (D) extending tablets in single-user [67] (Image © Hubenschmid et al., $C$ BY 4.0) and (E) multi-user scenarios [87] (Image © Langner et al., $C$ BY 4.0), and (F) finally shifting towards more empirical studies [70] (Image © Hubenschmid et al., $C$ BY 4.0). All images used with permission. To explore our investigated corpus, please visit https://imldresden.github.io/huis.
@@ 2.1 Background
The term HUI was coined in the early 1990s by Feiner and Shamash, describing a combination of “heterogeneous display and interaction device technologies” [47]. They argue that physical device sizes decreased with the advent of portable computing devices (i.e., laptops). This led to reduced interface real estate while retaining high-resolution input and output. In contrast, immersive technologies such as MR (especially HWDs)—in the early 1990s and still partly today—offer a lower resolution for both input and output space. However, AR has the potential for virtually unlimited interfaces that exceed the capabilities of conventional display technologies. Thus, Feiner and Shamash propose to combine these two technologies by “taking advantage of the strong points of each” [47], treating the technologies as complementary instead of competing. They exemplify this concept by presenting a “hybrid window manager”, combining a high-resolution yet restricted desktop interface with a low-resolution yet virtually unlimited AR head-mounted display interface into one unified application that blurs the boundaries between interfaces.
@@
This initially broad scope opened up a vast design space, but its ambiguity may have limited the adoption of the term within the field, leading to the diffusion of this term as it was applied to an increasing variety of device combinations over the years. Although one prominent theme was the combination of 2D devices and immersive MR environments, this was not universally shared by all prior works. On the one hand, Butz et al. highlighted that HUIs extend to various “technologies and techniques, including virtual elements such as 3D widgets, and physical objects such as tracked displays and input devices” [19]. They noted that the resulting global AR space can be shared, which is also discussed by Feiner, as HUIs combine all devices “in a mobile, shared environment” [45]. Bornik et al. further emphasize the potential combination of MR environments with conventional 2D display technology for HUIs, as they “pair 3D perception and direct 3D interaction with 2D system control and precise 2D interaction” [13]. This is echoed by Geiger et al., who state that HUIs “combine 2D, 3D, and real object interaction and may use multiple input and output devices and different modalities” [50]. In contrast, Strawhacker and Bers employ HUIs in a broader context without focusing on MR environments, presenting a HUI that allows “users [to] switch freely between tangible and graphical input” [138].
@@
One commonly shared theme is the importance of complementarity: For example, Sandor et al. state that “information [in HUIs] can be spread over a variety of different, but complementary, displays” [123]. Furthermore, in line with Butz et al. [19], they describe that users of HUIs can “interact through a wide range of interaction devices” [123]—we thus see the potential of HUIs not within a random assortment of technologies, but in a principled integration of different standalone “interaction devices”.
@@ 2.2 Adjacent Research Areas
The combination of complementary device technologies opens up a vast design space with countless possible combinations. These device combinations are, however, not exclusive to HUIs. In this section, we describe the intersection between HUIs and adjacent research and clearly position HUIs within the current research landscape.
@@ 2.2.1 Distributed User Interfaces
Elmqvist [38] defines the term distributed user interface as an interface whose components are distributed across one or more dimensions, such as input, output, platform, space, and time. This theoretical perspective can be seen as the foundation of any research on multi-device usage, including multimodal interaction. For HUIs, all proposed dimensions are relevant, as they span multiple technologies and distribute the user interface accordingly. However, by combining heterogeneous technologies, HUIs aim to distribute input and output to the most suitable device to achieve a goal while focusing on their complementary use. Therefore, we see distributed user interfaces as an overarching conceptual model, with HUIs representing one possible technical realization of this concept that focuses on heterogeneous devices.
@@ 2.2.2 Complementary Interfaces
Recent works by Zagermann et al. [154] as well as Elmqvist [39] highlight that attributing unique roles, properties, and purposes to individual devices and modalities can lead to meaningful combinations of interfaces that support users in their current task at hand. Such complementary interfaces distribute interaction across devices and modalities to establish a “symbiosis of interfaces, where each component purposefully increases the quality of interaction” [154]. In the context of HUIs, complementary interfaces can be considered as an overarching concept that includes combinations of homogeneous and heterogeneous devices, but also input (e.g., interaction techniques) and output modalities (e.g., visually or auditory). The core ideas of complementary interfaces are typically part of HUIs, in that HUIs represent a technical realization of this concept.
@@ 2.2.3 Cross-Device Interaction
Brudy et al. [16] provide a comprehensive overview of the field of cross-device interaction. Here, the focus is on research that “transcends the individual device and user” [16], which unifies research that is focused on different kinds of multi-device environments, ranging from multi-monitor setups to ad-hoc mobile device ecologies. Cross-device interaction is often related to Weiser’s vision of ubiquitous computing [149], studying devices such as smartphones, tablets, or larger interactive surfaces. Although Brudy et al. [16] list head-worn MR devices and tangibles as part of their cross-device taxonomy, they do not further elaborate on combinations of heterogeneous (e.g., non-immersive and immersive) devices.
@@ 2.2.4 Cross-Reality Environments
Using the reality–virtuality (RV) continuum of Milgram and Kishino [100] as a foundation, the research area of cross-reality environments investigates the benefits of combining various points on the RV continuum. Key aspects of cross-reality environments include “smooth transition[s] between systems using different degrees of virtuality” and “collaboration between users using different systems with different degrees of virtuality” [134]. Recent taxonomies by Auda et al. [4] and Wang and Maurer [147] further solidify concepts such as transitioning between different points on the RV continuum or concurrently using multiple distinct systems along the RV continuum.
For a single user, such cross-reality environments typically create a sequence of actions (e.g., switching devices). Depending on the workflow, this could be described as a migratory interface [16], asynchronous HUI [69], or transitional interface [98, 22]. Furthermore, given that previous work often focuses on HUIs that combine 2D interaction components with MR environments (e.g., [47, 13, 19, 50]), we see cross-reality as a potential umbrella term, where a subset of prior HUIs (see Sec. 2.3) lies at the intersection between cross-reality and cross-device interaction: While cross-device interaction focuses on the combination of devices, the area of cross-reality concentrates on the combination of realities. Thus, the area of cross-reality research encompasses both homogeneous (e.g., collaboration between HWDs in different realities) and heterogeneous device combinations (e.g., switching from desktop to VR). We consider the latter as part of our HUI terminology.
@@ 2.2.5 Transitional Interfaces
The MagicBook by Billinghurst et al. [11] is often described as the first transitional interface [98, 22]: Here, a user can move seamlessly along the RV continuum—from browsing the physical book to a handheld AR display to immersive VR. Although they can be regarded as a subset of cross-reality, transitional interfaces specifically focus on the design of transitions and their effect on users [98]. While early head-mounted displays required a discrete switch of hardware to move across interfaces, currently available HWDs allow for continuous transitions by interactively adding or removing virtual contents.
Similarly to research on cross-reality environments, transitional interfaces typically involve a sequence of activities: A user solves one aspect of a given task in AR and continues to work on the activity in VR. The transition keeps the user oriented in the task space, choosing one manifestation of the RV continuum at a time. However, for HUIs, connecting to multiple manifestations at the same time is a key ability. In line with previous research, we “think of HUIs and transitional interfaces as complementary” [22], rather than competing: A HUI may use a transitional interface when switching between devices; likewise, a transitional interface may encompass multiple heterogeneous devices to be switched between.
@@ 2.2.6 Further Related Terms
Adjacent research areas additionally use related terms to describe their work: The term cross-device interaction is often used interchangeably with the term “cross-surface interaction”, while earlier work in that area is often considered as “multi-device” or “multi-display” systems [16]. Similarly, there are other terms that are related to HUIs, but focus on interaction techniques (i.e., “hybrid interaction” [81, 9]) or the setting (i.e., “hybrid virtual environment” [22, 32, 145]). The term “augmented displays” was previously used to describe integral concepts of HUIs (cf. [118, 120]). Therefore, we consider it as a unique configuration for HUIs (see also Sec. 4).
2.3 Identifying Hybrid User Interfaces

### v2

@@ Abstract
We investigate hybrid user interfaces (HUIs), aiming to establish a cohesive understanding and to adopt consistent terminology for this nascent research area. HUIs combine heterogeneous devices in complementary roles, leveraging the distinct benefits of each. Our work focuses on cross-device interaction between 2D devices and mixed reality environments, which are particularly compelling, leveraging the familiarity of traditional 2D platforms while providing spatial awareness and immersion. Although prior work has prominently explored such HUIs in the context of mixed reality, we still lack a cohesive understanding of the unique design possibilities and challenges of such combinations, resulting in a fragmented research landscape. We conducted a systematic survey and present a taxonomy of HUIs that combine conventional display technology and mixed reality environments. Based on this, we discuss past and current challenges, the evolution of definitions, and prospective opportunities to tie together the past 30 years of research with our vision of future HUIs.
@@ 1 Introduction
Immersive augmented reality (AR) and virtual reality (VR) is gradually gaining relevance in everyday life, with affordable off-the-shelf hardware becoming increasingly available to consumers. These mixed reality (MR) platforms [^1] In this work, mixed reality refers to both AR and VR (cf. [speicher2019what]), as discussed in Sec. 4.3. have now evolved to the point where researchers can explore the nuances of interaction design without being constrained by major technological limitations. In this context, Feiner and Shamash [feiner1991hybrid] proposed in 1991 the concept of hybrid user interfaces (HUIs), which combine “heterogeneous display and interaction device technologies.” This concept theoretically allows for an infinite integration of technologies, yet, such heterogeneous combinations are especially compelling for MR environments: The ubiquity, convenience, and familiarity of conventional 2D platforms (e.g., smartphones, desktops) provide a perfect complement to the immersion and complexity of optical see-through (OST) and video see-through (VST) head-worn devices (HWDs) [^2] We refer to “head-worn devices” (HWDs) instead of “head-mounted displays” (HMDs) to emphasize the increase in wearability and capabilities of current hardware, but intentionally kept the previous term HMD for describing older hardware.. Yet, although the concept of HUIs has persisted for decades, no coherent delineation has emerged. As a result, there is a distinct lack of consistent design models and terminologies—fragmenting the research community across overlapping research areas such as cross-reality systems [auda2023scoping], transitional interfaces [grasset2006transitional], or cross-device interaction [brudy2019crossdevice].
For example, the seminal cross-device taxonomy by Brudy et al. [brudy2019crossdevice] provides an overarching model for the research area of HUIs. However, their taxonomy only considers established devices (e.g., mobile devices) and homogeneous combinations, leaving topics such as MR environments largely for future work. Yet, without firm design principles and in the face of countless possible device combinations, the design space of HUIs can appear bewilderingly large. Therefore, we extend the existing taxonomy by examining complementary cross-device interaction between traditional 2D device technologies and novel MR platforms as a prevalent subset of HUIs. By first examining this subset, we take a first step towards a better understanding of the broader research area of HUIs. Thus, we aim to establish a common terminology, allowing researchers and practitioners to better benefit from shared insights, establish a consistent framework, and inspire future systems. This would, in turn, enable the creation of a cohesive understanding of the unique design possibilities and challenges of HUIs.
@@
  • A positioning of MR-HUIs, creating a better understanding of the term within the current fragmented research landscape (Sec. 2 and Sec. 7).
@@ 2 What is a Hybrid User Interface?
Since 1991, the term “hybrid user interface” has continued to evolve. Despite significant advances in device technologies, no coherent definition has emerged that delineates HUIs from related research areas. This lack of common understanding makes it difficult to compare results and share insights. In the following, we first review prior definitions and the background (Sec. 2.1) of HUIs, building on previous work by Satkowski and Méndez [satkowski2023fantastic], and position HUIs in relation to current adjacent research areas (Sec. 2.2). Based on this, we discuss how we can identify HUIs (Sec. 2.3) and derive three attributes that specifically characterize HUIs that combine MR environments with conventional display technologies (Sec. 2.4) to guide our literature survey.
Figure 1: We take a look at 30 years of hybrid user interface research and its evolution, focusing on combinations of conventional 2D devices with mixed reality environments. From left to right, (A) starting with its definition [feiner1991hybrid], (B) early extensions [bornik2006hybrid] (Image © IEEE 2006), and (C) current usage for augmenting displays [reipschlager2021personal] (Image © Reipschläger et al., $C$ BY 4.0), (D) extending tablets in single-user [hubenschmid2021stream] (Image © Hubenschmid et al., $C$ BY 4.0) and (E) multi-user scenarios [langner2021marvis] (Image © Langner et al., $C$ BY 4.0), and (F) finally shifting towards more empirical studies [hubenschmid2023smartphone] (Image © Hubenschmid et al., $C$ BY 4.0). All images used with permission. To explore our investigated corpus, please visit https://imldresden.github.io/huis.
@@ 2.1 Background
The term HUI was coined in the early 1990s by Feiner and Shamash, describing a combination of “heterogeneous display and interaction device technologies” [feiner1991hybrid]. They argue that physical device sizes decreased with the advent of portable computing devices (i.e., laptops). This led to reduced interface real estate while still retaining high-resolution input and output. In contrast, immersive technologies such as MR (especially HWDs)—back then and still mostly today—offer a lower resolution for both input and output space. However, AR has the potential for virtually unlimited interfaces that exceed the capabilities of conventional display technologies. Thus, they propose to combine these technologies by “taking advantage of the strong points of each” [feiner1991hybrid], treating the technologies as complementary instead of competing. They exemplify this concept in a “hybrid window manager”, combining a high-resolution yet size-restricted desktop with a low-resolution yet virtually unlimited AR head-mounted display interface into a unified application that blurs the boundaries between interfaces.
@@
This initially broad scope opened up a vast design space, but its ambiguity may have limited the adoption of the term within the field, leading to the diffusion of this term as it was applied to an increasing variety of device combinations over the years. Although one prominent theme was the combination of 2D devices and immersive MR environments, this was not universally shared by all prior works. On the one hand, Butz et al. highlighted that HUIs extend to various “technologies and techniques, including virtual elements such as 3D widgets, and physical objects such as tracked displays and input devices” [butz1999enveloping]. They noted that the resulting global AR space can be shared, which is also discussed by Feiner, as HUIs combine all devices “in a mobile, shared environment” [feiner2000environment]. Bornik et al. further emphasize the potential combination of MR environments with conventional 2D display technology for HUIs, as they “pair 3D perception and direct 3D interaction with 2D system control and precise 2D interaction” [bornik2006hybrid]. This is echoed by Geiger et al., who state that HUIs “combine 2D, 3D, and real object interaction and may use multiple input and output devices and different modalities” [geiger2008hyui]. In contrast, Strawhacker and Bers employ HUIs in a broader context without MR, presenting a HUI that allows “users [to] switch freely between tangible and graphical input” [strawhacker2015want].
@@
One commonly shared theme is the importance of complementarity: For example, Sandor et al. state that “information [in HUIs] can be spread over a variety of different, but complementary, displays” [sandor2005immersive]. Furthermore, in line with Butz et al. [butz1999enveloping], they describe that users of HUIs can “interact through a wide range of interaction devices” [sandor2005immersive]—we thus see the potential of HUIs not within a random assortment of technologies, but in a principled integration of different standalone “interaction devices”.
TABLE I: An overview of adjacent research areas and their relation to HUIs.
Research Area Description Relation to HUIs
Distributed User Interfaces [elmqvist2011distributed] Foundation for all research on multi-device usage. HUIs are one possible realization, focusing on distributing interfaces across heterogeneous devices.
Complementary Interfaces [zagermann2022complementary] Symbiotic combinations of devices and interfaces through complementarity. HUIs focus on complementarity between devices and thus are a technological approach to complementary interfaces.
Cross-Device Interaction [brudy2019crossdevice] Describes general research on homogeneous and heterogeneous device combinations. HUIs are a subset of cross-device interaction focusing on heterogeneous combinations.
Cross-Reality Environments [auda2023scoping] Systems and interfaces that span across different points on the reality-virtuality continuum. Significant amounts of prior work in HUIs combine MR environments (e.g., AR and VR) with conventional displays (i.e., reality), thus representing a subset of cross-reality.
Transitional Interfaces [mayer2024crossing] Describes transitions between environments, interfaces, and devices. HUIs may use transitional interfaces when switching between devices; likewise, transitional interfaces may encompass multiple heterogeneous devices to be switched between.
@@ 2.2 Adjacent Research Areas
The combination of complementary device technologies opens up a vast design space with countless possible combinations. These device combinations are, however, not exclusive to HUIs. In this section, we describe the intersection between HUIs and adjacent research and clearly position HUIs within the current research landscape (see also Tab. I for an overview).
@@ 2.2.1 Distributed User Interfaces
Elmqvist [elmqvist2011distributed] defines the term distributed user interface as an interface whose components are distributed across one or more dimensions, such as input, output, platform, space, and time. This theoretical perspective can be seen as the foundation of any research on multi-device usage, including multimodal interaction. For HUIs, all proposed dimensions are relevant, as they span multiple technologies and distribute the user interface accordingly. However, by combining heterogeneous technologies, HUIs aim to distribute input and output to the most suitable device to achieve a goal while focusing on their complementary use. Therefore, we see distributed user interfaces as an overarching conceptual model, with HUIs representing a possible technical realization of this concept that focuses on heterogeneous devices.
@@ 2.2.2 Complementary Interfaces
Recent works by Zagermann et al. [zagermann2022complementary] as well as Elmqvist [elmqvist2023anywhere] highlight that attributing unique roles, properties, and purposes to individual devices and modalities can lead to meaningful combinations of interfaces that support users in their current task at hand. Such complementary interfaces distribute interaction across devices and modalities to establish a “symbiosis of interfaces, where each component purposefully increases the quality of interaction” [zagermann2022complementary]. In the context of HUIs, complementary interfaces can be considered as an overarching concept that includes combinations of homogeneous and heterogeneous devices, but also input (e.g., interaction techniques) and output modalities (e.g., visually or auditory). The core ideas of complementary interfaces are typically part of HUIs, in that HUIs represent a technical realization of this concept.
@@ 2.2.3 Cross-Device Interaction
Brudy et al. [brudy2019crossdevice] provide a comprehensive overview of the field of cross-device interaction. Here, the focus is on research that “transcends the individual device and user” [brudy2019crossdevice], which unifies research that is focused on different kinds of multi-device environments, ranging from multi-monitor setups to ad-hoc mobile device ecologies. Although Brudy et al. [brudy2019crossdevice] list head-worn MR devices and tangibles as part of their cross-device taxonomy, they do not further elaborate on combinations of heterogeneous (e.g., non-immersive and immersive) devices.
@@ 2.2.4 Cross-Reality Environments
Using the reality–virtuality (RV) continuum [milgram1994taxonomy] as a foundation, the research area of cross-reality environments investigates the benefits of combining various points on the RV continuum. Key aspects of cross-reality environments include “smooth transition[s] between systems using different degrees of virtuality” and “collaboration between users using different systems with different degrees of virtuality” [simeone2020international]. Recent taxonomies by Auda et al. [auda2023scoping] and Wang and Maurer [wang2022design] further solidify concepts such as transitioning between different points on the RV continuum or concurrently using multiple distinct systems along the RV continuum.
For a single user, such cross-reality environments typically create a sequence of actions (e.g., switching devices). Depending on the workflow, this could be described as a migratory interface [brudy2019crossdevice], asynchronous HUI [hubenschmid2021asynchronous], or transitional interface [mayer2024crossing, carvalho2012design]. Furthermore, given that previous work often focuses on HUIs that combine 2D IO components with MR environments (e.g., [feiner1991hybrid, bornik2006hybrid, butz1999enveloping, geiger2008hyui]), we see cross-reality as a potential umbrella term, where a subset of prior HUIs (see Sec. 2.3) lies at the intersection between cross-reality and cross-device interaction: While cross-device interaction focuses on the combination of devices, the area of cross-reality concentrates on the combination of realities. Thus, the area of cross-reality research encompasses both homogeneous (e.g., collaboration between HWDs in different realities) and heterogeneous device combinations (e.g., switching from desktop to VR). We consider the latter as part of our HUI terminology.
@@ 2.2.5 Transitional Interfaces
The MagicBook by Billinghurst et al. [billinghurst2001magicbook] is often described as the first transitional interface [mayer2024crossing, carvalho2012design]: Here, a user can move seamlessly along the RV continuum—from browsing the physical book to a handheld AR display to immersive VR. Although they can be regarded as a subset of cross-reality, transitional interfaces specifically focus on the design of transitions and their effect on users [mayer2024crossing]. While early head-mounted displays required a discrete switch of hardware to move across interfaces, currently available HWDs allow for continuous transitions by interactively adding or removing virtual contents.
Similarly to research on cross-reality environments, transitional interfaces typically involve a sequence of activities: A user solves one aspect of a given task in AR and continues to work on the activity in VR. The transition keeps the user oriented in the task space, choosing one manifestation of the RV continuum at a time. However, for HUIs, connecting to multiple manifestations at the same time is a key ability. In line with previous research, we “think of HUIs and transitional interfaces as complementary” [carvalho2012design], rather than competing: A HUI may use a transitional interface when switching between devices; likewise, a transitional interface may encompass multiple heterogeneous devices to be switched between.
@@ 2.2.6 Further Related Terms
Adjacent research areas additionally use related terms to describe their work: The term cross-device interaction is often used interchangeably with the term “cross-surface interaction”, while earlier work in that area is often considered as “multi-device” or “multi-display” systems [brudy2019crossdevice]. Similarly, there are other terms that are related to HUIs, but focus on interaction techniques (i.e., “hybrid interaction” [knierim2021smartphone, besancon2021state]) or the setting (i.e., “hybrid virtual environment” [carvalho2012design, decarvalho2009designing, wang2014coordinated]). The term “augmented displays” was previously used to describe integral concepts of HUIs (cf. [reipschlager2020augmented, reipschlager2018debugar]). Therefore, we consider it as a unique configuration for HUIs.
2.3 Characterizing Hybrid User Interfaces